A magnetic B field of strength 4 T is perpendicular to a conducting loop of wire measuring 0.4 square meters with 50 turns. If the B field is linearly reduced to zero in 2 s, then what is the magnitude of the induced emf voltage? (in units of V)
Answer #1

Magnetic flux,\large \phi=BAN

where B is the magnetic field,A is the area of the loop and N is the number of turns.

Here,

Initial magnetic field,\large B_0= 4 T

Area of the loop = 0.4 \large m^2

number of turns, N = 50

Final magnetic field,\large B_t=0

Time , t = 2s

Now,

Initial magnetic flux,\large \phi_1 = B_0AN =4\times 0.4\times 50 = 80 Wb

Final magnetic flux,\large \phi_2 = B_tAN =0\times 0.4\times 50 = 0

Induced voltage = \large \frac{\phi_1-\phi_2}{t}=\frac{80-0}{2}=40V